OpenStudy (anonymous):

Jackson ordered some clothes for school online. He bought 2 pairs of pants and 5 shirts. He spent a total of $138.45. Part 1: Write the equation for this word problem. Be sure to identify your variables. Part 2: If each pair of pants cost $14.25, solve the equation to find the cost of each shirt. Show your work.

OpenStudy (praxer):

2x+5y =$138.45 x is the price of 1pair of pant y is the price of 1 pair of shirt now solve the equation by putting the value $14.25 in the next

OpenStudy (praxer):

$14.25*2+5y=$138.45 5y= $138.5-$28.5 y=$110/5 y=$22

OpenStudy (anonymous):

thanks can you help me on one more please